# TFusionOcc: Student’s t-Distribution Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction

### v1

@@ IV-G Performance Analysis On nuScenes-C Dataset
Method Metric Clean Cam Crash Frame Lost Color Quant Motion Blur Bright Low Light Fog Snow OccCylindrical IoU 44.94 41.76 (-3.18%) 40.84 (-4.10%) 43.64 (-1.30%) 41.45 (-3.49%) 44.78 (-0.16%) 43.48 (-1.46%) 44.59 (-0.35%) 38.92 (-6.02%) DAOcc IoU 45.00 42.75 (-2.25%) 42.16 (-2.84%) 43.88 (-1.12%) 42.75 (-2.25%) 43.72 (-1.28%) 41.25 (-3.75%) 43.92 (-1.08%) 40.72 (-4.28%) TFusionOcc (T-P-25600) IoU 45.49 42.59 (-2.90%) 41.97 (-3.52%) 43.14 (-2.35%) 40.45 (-5.04%) 44.81 (-0.68%) 41.24 (-4.25%) 44.14 (-1.35%) 40.58 (-4.91%) TFusionOcc (T-SQ-25600) IoU 46.14 42.59 (-3.55%) 41.93 (-4.21%) 43.03 (-3.11%) 38.36 (-7.78%) 45.41 (-0.73%) 41.99 (-4.15%) 44.90 (-1.24%) 40.09 (-6.05%) TFusionOcc (T-SQ-IW-25600) IoU 47.01 42.61 (-4.40%) 41.77 (-5.24%) 43.97 (-3.04%) 38.58 (-8.43%) 46.31 (-0.70%) 42.19 (-4.82%) 45.95 (-1.06%) 40.37 (-6.64%)
TABLE VI: 3D semantic occupancy prediction IoU results on nuScenes-C Camera Corruption benchmark. All methods are evaluated with different corruption type input images provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
Method Metric Clean Cam Crash Frame Lost Color Quant Motion Blur Bright Low Light Fog Snow SurroundOcc mIoU 20.30 11.60 (-8.70%) 10.00 (-10.30%) 14.03 (-6.27%) 12.41 (-7.89%) 19.18 (-1.12%) 12.15 (-8.15%) 18.42 (-1.88%) 7.39 (-12.91%) Inverse++ mIoU 20.91 11.76 (-9.15%) 10.22 (-10.69%) 14.88 (-6.03%) 12.84 (-8.07%) 19.71 (-1.20%) 12.05 (-8.86%) 18.86 (-2.05%) 7.62 (-13.29%) OccCylindrical mIoU 28.67 23.38 (-5.29%) 21.87 (-6.80%) 25.38 (-3.29%) 22.57 (-6.10%) 27.88 (-0.79%) 22.17 (-6.50%) 27.38 (-1.29%) 18.29 (-10.38%) DAOcc mIoU 30.50 22.35 (-8.15%) 20.78 (-9.72%) 27.46 (-3.04%) 26.25 (-4.25%) 26.78 (-3.72%) 21.15 (-9.35%) 26.87 (-3.63%) 19.42 (-11.08%) TFusionOcc (T-P-25600) mIoU 30.04 22.34 (-7.70%) 20.78 (-9.26%) 24.42 (-5.62%) 20.04 (-10.00%) 29.03 (-1.01%) 18.62 (-11.42%) 27.43 (-2.61%) 15.62 (-14.42%) TFusionOcc (T-SQ-25600) mIoU 30.88 23.50 (-7.38%) 22.04 (-8.84%) 25.65 (-5.23%) 20.25 (-10.63%) 29.92 (-0.96%) 20.98 (-9.90%) 28.85 (-2.03%) 16.33 (-14.55%) TFusionOcc (T-SQ-IW-25600) mIoU 31.47 24.39 (-7.08%) 23.03 (-8.44%) 26.14 (-5.33%) 21.06 (-10.41%) 30.58 (-0.89%) 21.67 (-9.80%) 29.69 (-1.78%) 17.45 (-14.02%)
TABLE VII: 3D semantic occupancy prediction mIoU results on nuScenes-C Camera Corruption benchmark. All methods are evaluated with different corruption type input images provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
Method Metric Clean Beam Missing Cross Sensor Cross Talk Incomplete Echo Motion Blur Wet Ground Fog Snow OccCylindrical IoU 35.89 30.73 (-5.16%) 25.01 (-10.88%) 33.29 (-2.60%) 35.37 (-0.52%) 32.44 (-3.45%) 35.04 (-0.85%) 25.79 (-10.10%) 34.00 (-1.89%) DAOcc IoU 42.39 38.17 (-4.22%) 33.79 (-8.60%) 33.68 (-8.71%) 42.01 (-0.38%) 38.05 (-4.34%) 41.67 (-0.72%) 31.47 (-10.92%) 37.01 (-5.38%) TFusionOcc (T-P-25600) IoU 40.29 32.75 (-7.54%) 25.65 (-14.64%) 36.25 (-4.04%) 39.67 (-0.62%) 34.79 (-5.50%) 38.17 (-2.12%) 26.44 (-13.85%) 36.88 (-3.41%) TFusionOcc (T-SQ-25600) IoU 40.53 33.00 (-7.53%) 25.89 (-14.64%) 36.60 (-3.93%) 39.90 (-0.63%) 34.67 (-5.86%) 38.74 (-1.79%) 26.54 (-13.99%) 36.88 (-3.65%) TFusionOcc (T-SQ-IW-25600) IoU 41.53 34.09 (-7.44%) 26.82 (-14.71%) 37.50 (-4.03%) 40.94 (-0.59%) 35.54 (-5.99%) 39.84 (-1.69%) 27.10 (-14.43%) 37.70 (-3.83%)
TABLE VIII: 3D semantic occupancy prediction IoU results on nuScenes-C Lidar Corruption benchmark. All methods are evaluated with different corruption type input images provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
Method Metric Clean Beam Missing Cross Sensor Cross Talk Incomplete Echo Motion Blur Wet Ground Fog Snow OccCylindrical mIoU 24.06 19.83 (-4.23%) 16.20 (-7.86%) 22.60 (-1.46%) 20.93 (-3.13%) 20.85 (-3.21%) 23.31 (-0.75%) 17.32 (-6.74%) 22.38 (-1.68%) DAOcc mIoU 27.73 22.96 (-4.77%) 18.41 (-9.32%) 22.86 (-4.87%) 24.45 (-3.28%) 23.66 (-4.07%) 27.25 (-0.48%) 19.73 (-8.00%) 24.53 (-3.20%) TFusionOcc (T-P-25600) mIoU 26.13 20.49 (-5.64%) 15.38 (-10.75%) 23.76 (-2.37%) 22.69 (-3.44%) 21.39 (-4.74%) 25.05 (-1.08%) 17.44 (-8.69%) 23.15 (-2.98%) TFusionOcc (T-SQ-25600) mIoU 27.03 21.07 (-5.96%) 15.61 (-11.42%) 24.73 (-2.30%) 23.30 (-3.73%) 22.15 (-4.88%) 26.06 (-0.97%) 18.40 (-8.63%) 24.37 (-2.66%) TFusionOcc (T-SQ-IW-25600) mIoU 27.50 21.81 (-5.69%) 16.49 (-11.01%) 24.94 (-2.56%) 24.11 (-3.39%) 22.62 (-4.88%) 26.57 (-0.93%) 18.63 (-8.87%) 24.37 (-3.13%)
TABLE IX: 3D semantic occupancy prediction mIoU results on nuScenes-C Lidar Corruption benchmark. All methods are evaluated with different corruption type input lidar point cloud provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
Corruption Severity IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation CameraCrash Easy 43.16 26.44 23.34 16.62 28.80 35.79 19.34 21.30 28.47 19.01 18.46 24.14 36.89 19.19 25.07 27.42 35.32 43.94 CameraCrash Mid 42.27 23.37 17.66 11.56 26.79 33.91 14.58 16.88 26.48 16.44 13.69 19.75 35.87 16.54 22.43 24.70 34.07 42.56 CameraCrash Hard 42.39 23.35 20.30 11.23 26.54 34.40 15.27 13.93 25.80 17.87 11.15 21.39 35.24 14.59 22.44 25.18 34.52 43.77 FrameLost Easy 44.33 27.42 25.40 16.12 30.44 36.89 19.15 21.10 28.59 20.72 18.97 26.33 38.65 21.48 26.20 28.11 35.81 44.84 FrameLost Mid 41.33 22.19 15.55 10.88 24.67 33.26 13.95 14.55 25.62 15.66 11.70 19.94 33.63 13.21 21.59 24.17 33.98 42.73 FrameLost Hard 39.66 19.48 9.91 9.09 22.02 31.26 11.74 11.63 23.98 12.70 8.77 15.30 30.99 9.11 18.80 21.75 33.03 41.52 ColorQuant Easy 46.55 30.61 31.18 19.09 33.46 39.63 22.86 26.15 30.57 23.20 22.87 30.29 42.26 25.40 28.71 30.42 37.19 46.44 ColorQuant Mid 44.19 26.96 25.94 14.03 30.21 38.26 19.43 18.13 29.03 22.00 18.56 26.63 38.54 19.68 24.39 26.26 35.72 44.54 ColorQuant Hard 41.18 20.85 13.87 8.24 22.61 34.82 12.93 8.22 26.70 18.66 9.86 20.29 33.70 9.37 19.14 33.79 42.30 20.85 MotionBlur Easy 41.43 26.89 28.45 15.39 31.42 37.33 20.59 19.98 29.74 23.25 19.18 25.99 36.16 17.03 22.63 25.86 34.73 42.48 MotionBlur Mid 37.26 18.80 15.08 8.51 21.43 33.12 11.83 13.43 22.58 17.93 6.36 13.80 29.91 3.61 12.99 20.31 31.56 38.34 MotionBlur Hard 37.06 17.48 12.18 7.54 19.65 32.08 9.80 12.42 20.60 16.37 4.50 12.63 29.47 2.63 11.26 19.46 31.17 37.93 Brightness Easy 46.90 31.35 31.83 20.44 34.29 39.73 23.55 27.70 31.19 23.89 23.07 30.52 43.06 27.37 29.94 30.92 37.45 46.57 Brightness Mid 46.45 30.67 31.12 19.86 32.69 39.37 22.29 26.38 31.08 23.73 22.70 29.40 42.35 27.07 28.96 30.20 37.26 46.31 Brightness Hard 45.59 29.71 30.29 19.08 31.22 38.80 20.78 25.06 30.92 23.76 21.97 28.14 40.90 25.49 27.29 28.93 36.92 45.88 LowLight Easy 43.86 25.38 19.52 14.11 30.88 37.84 16.55 20.19 28.43 17.38 13.94 25.64 37.39 15.11 23.22 25.82 35.57 44.48 LowLight Mid 42.12 21.69 11.25 11.54 27.65 36.35 11.88 18.08 26.75 12.07 5.65 22.09 34.43 8.43 19.90 23.02 34.45 43.44 LowLight Hard 40.60 17.93 3.52 6.49 23.37 34.05 7.87 15.49 23.81 5.33 1.57 17.02 32.01 3.56 16.69 20.34 33.18 42.59 Fog Easy 46.37 30.57 30.82 19.55 34.30 39.40 21.91 25.99 30.65 23.35 22.12 30.06 42.13 26.77 28.91 29.86 37.10 46.27 Fog Mid 46.03 29.89 30.01 18.93 33.72 39.29 20.89 25.12 30.42 22.91 21.21 29.32 41.37 25.36 27.89 28.93 36.87 46.07 Fog Hard 45.44 28.61 29.00 17.43 32.02 38.90 18.92 24.15 30.04 22.41 19.61 27.90 40.02 22.29 25.69 27.37 36.47 45.58 Snow Easy 41.26 20.46 12.34 8.18 24.01 34.48 14.53 13.38 26.04 20.12 6.13 19.37 32.52 6.58 12.73 21.68 33.02 42.31 Snow Mid 39.61 15.81 7.13 4.17 16.34 30.79 7.63 11.86 22.00 16.87 4.93 9.97 27.63 2.30 7.69 15.13 30.04 38.46 Snow Hard 40.23 16.08 5.29 4.19 16.00 31.17 8.68 11.28 22.39 16.59 4.58 10.08 29.80 2.64 9.17 14.27 31.40 39.78
TABLE X: The performance of TFusionOcc(T-SQ-IW-25600) on the nuScene-C [31] dataset under different camera corruption and severity settings.
Corruption Severity IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation BeamMissing Light 38.03 24.83 25.39 12.35 30.38 33.20 17.80 20.84 23.97 16.98 19.35 26.19 37.83 21.87 23.66 24.77 26.92 35.79 BeamMissing Moderate 34.28 21.94 22.45 9.84 28.26 29.26 15.39 17.79 20.83 12.97 17.61 23.96 35.44 19.69 20.99 22.06 23.07 31.41 BeamMissing Heavy 29.97 18.67 18.85 6.93 25.51 24.95 13.02 13.23 15.80 11.40 15.58 21.18 32.63 17.35 18.16 18.45 19.03 26.71 CrossSensor Light 31.71 20.08 19.55 7.95 27.85 27.18 14.63 13.75 17.86 11.99 17.34 23.06 34.50 19.05 19.57 19.78 20.08 27.19 CrossSensor Moderate 26.52 16.07 15.50 5.00 23.42 21.71 12.15 7.42 13.75 9.30 14.33 19.55 31.52 15.85 16.60 16.43 14.97 19.66 CrossSensor Heavy 22.23 13.32 11.50 3.77 21.31 16.56 6.96 9.04 10.43 7.01 13.48 15.76 28.99 14.64 14.00 13.66 11.28 14.64 CrossTalk Light 39.77 26.33 28.50 14.08 31.14 35.44 18.51 21.00 25.73 17.29 19.97 27.30 39.52 24.07 26.08 27.46 29.48 35.63 CrossTalk Moderate 37.47 24.93 27.82 12.62 30.10 34.11 16.63 19.01 24.07 15.91 19.14 26.14 38.85 23.78 25.07 26.82 28.03 30.82 CrossTalk Heavy 35.26 23.57 26.75 11.55 29.19 32.80 15.39 16.61 22.60 14.60 18.47 25.09 38.09 23.34 23.41 25.86 26.44 26.93 IncompleteEcho Light 41.18 25.74 28.97 12.08 28.39 30.45 15.70 19.21 27.68 19.35 17.63 24.15 39.80 23.88 26.37 27.75 30.51 39.94 IncompleteEcho Moderate 41.01 24.66 28.96 10.54 25.72 26.81 13.15 17.27 27.67 19.26 15.40 21.58 39.74 23.90 26.36 27.73 30.52 39.93 IncompleteEcho Heavy 40.62 21.93 28.92 8.66 17.25 16.89 7.41 13.52 27.62 19.08 8.97 14.64 39.57 23.92 26.34 27.71 30.53 39.91 MotionBlur Light 38.10 24.89 25.61 12.99 30.89 33.74 19.39 20.91 24.15 14.99 19.40 26.77 36.30 21.25 23.10 24.59 26.09 38.00 MotionBlur Moderate 35.44 22.56 22.40 11.01 29.50 31.09 18.24 18.10 20.87 11.53 18.13 25.43 33.88 18.98 20.68 22.26 22.74 36.10 MotionBlur Heavy 33.07 20.42 19.24 9.15 28.01 28.37 17.26 15.57 17.81 8.80 16.49 23.64 32.07 17.44 18.77 20.20 19.85 34.03 WetGround Light 40.53 26.99 28.86 15.18 31.66 36.33 19.80 23.18 27.60 19.65 20.73 28.03 37.07 21.62 23.82 27.84 30.50 39.99 WetGround Moderate 39.84 26.57 28.86 15.12 31.65 36.23 19.77 22.70 27.53 19.86 20.71 27.91 35.13 19.21 21.98 28.01 30.50 39.98 WetGround Heavy 39.14 26.14 28.80 14.79 31.57 36.14 19.74 22.69 27.47 19.87 20.68 27.85 33.08 17.10 20.01 27.99 30.49 39.98 Fog Light 36.86 24.38 27.36 14.09 28.96 33.78 15.06 20.04 23.85 18.70 17.81 23.49 38.44 22.32 24.71 23.95 25.84 31.69 Fog Moderate 28.39 20.05 22.93 12.52 26.29 29.36 12.23 17.41 19.34 16.90 11.32 16.80 34.83 20.18 21.50 19.90 16.92 22.37 Fog Heavy 16.04 11.46 13.35 6.10 15.92 19.10 7.93 8.75 9.78 10.17 3.34 7.81 26.48 14.66 12.86 11.56 6.17 9.35 Snow Light 39.38 24.78 28.29 14.52 29.78 34.99 14.71 19.91 23.43 17.00 18.80 25.34 39.33 23.29 25.23 26.10 27.60 28.23 Snow Moderate 37.52 24.42 28.34 14.26 29.77 34.75 14.91 19.96 22.36 15.96 18.51 24.98 38.93 23.35 25.34 26.44 27.15 25.73 Snow Heavy 36.20 23.91 28.11 13.47 29.74 34.20 15.34 18.98 21.07 13.80 18.17 25.16 38.42 23.08 25.28 26.47 25.68 25.65
TABLE XI: The performance of TFusionOcc(T-SQ-IW-25600) on the nuScene-C [31] dataset under different lidar corruption and severity settings.
IV-H Performance Analysis On Varying Distance
IV-H1 Performance Variation Trend At Different Distances.
Following [18], we take the ego vehicle as the origin and $R$ as the radius. By adjusting the length of $R$, we evaluate our proposed method against other SOTA algorithms at varying perception distance radius [20m, 25m, 30m, 35m, 40m, 45m, 50m], including challenging rainy and night scenarios. The trend of performance variation is shown in Figure 10.
The trend of performance variation across the whole validation set, as shown in Figures 10(a) and 10(d), aligns with the performance benchmark in Table II. In the rainy scenario, the performance variation trend of our proposed method is similar to the results of the whole validation set experiment, except that the performance gap between different T-Primitive settings further narrows down to 0.11% to 2.90% under the mIoU metric and 0.44% to 1.74% under the IoU metric due to the semantic and geometric information loss from both sensors. In the nighttime scenario, we discover that our T-SQ-25600 setting outperforms the second-best T-SQ-IW-12800 setting within a 35-meter perception range under the mIoU metric, as shown in Figure 10(c). This phenomenon may stem from the fact that more primitives can better capture close-range details, leading to better robustness against challenging dim-lighting conditions.
Figure 10: Performance variation trend for 3D semantic occupancy prediction task. (a) mIoU performance variation trend on the whole SurroundOcc-nuScenes validation set, (b) IoU performance variation on the whole SurroundOcc-nuScenes validation set, (c) mIoU performance variation trend on the SurroundOcc-nuScenes validation rainy scenario subset, (d) IoU performance variation on the SurroundOcc-nuScenes validation rainy scenario subset, (e) mIoU performance variation on the SurroundOcc-nuScenes validation night scenario subset, and (f) IoU performance variation on the SurroundOcc-nuScenes validation night scenario subset. Better viewed when zoomed in.
IV-H2 Performance In Different Sectors Ranges
[…]
@@ IV-H2 Performance In Different Sectors Ranges
IV-I Qualitative Study
IV-I1 Qualitative Results On SurroundOcc-nuScenes Dataset
To qualitatively study the difference in each setting of the proposed TFusionOcc, we demonstrate the visualization result of each setting in daytime, rainy, and night-time scenarios in Figure 12, with the main discrepancy area highlighted. In the daytime scenarios as shown at the top of Figure 12, in the orange square highlighted area, except for the T-P-25600 setting, all other settings successfully detect those two pedestrians. What is more, T-SQ-IW Primitive preserves better the geometric contour shape of background buildings, such as sharper building edges and more accurate building boundaries, than T-P primitive and T-SQ primitive, demonstrating the importance of the geometric shape flexibility of intermediate primitives. In rainy scenarios, as shown in the middle of Figure 12, in the dark red square highlighted area. The better geometric shape representation capability of the primitives could more accurately capture the geometric structure of the object, leading to better detection performance; T-SQ-IW-25600 achieved the best performance due to its best geometric shape representation capability, and T-P-25600 the worst, stemming from the ellipsoid geometric shape prior. Similarly, in the nighttime scenarios, as shown in the bottom of Figure 12, in the square highlighted area. Even with severe occlusions and a vehicle headlight glare pollution problem, we observed that greater geometric shape flexibility of primitives leads to better detection performance, which aligns with rainy and daytime phenomena.
Figure 12: Qualitative results under different settings for daytime, rainy, and nighttime scenarios displayed in the upper, middle, and bottom sections, respectively. Better viewed when zoomed in.
Furthermore, to demonstrate the effectiveness and robustness of our method, we compare TFusionOcc in three different T-Primitive settings, such as T-P-25600, T-SQ-25600 and T-SQ-IW-25600, against other SOTA algorithms in challenging daytime, rainy and nighttime scenarios, as shown in Figure 13, with the primary difference area highlighted in each scenario. In the daytime scenario, as shown at the top of Figure 13, in the orange square highlighted area, benefit from multi-task learning, which provides an additional supervision signal, DAOcc successfully detects distant pedestrians and achieves similar performance as TFusionOcc. However, in the dark-red-square-highlighted area, DAOcc fails to detect the distant vehicle, demonstrating the superior performance of our proposed method. In the rainy scenario, as shown in the middle of Figure 13, in the purple square highlighted area, our proposed method not only successfully detects all dynamic objects, but also preserves their general contours well. In the highlighted dark-blue-square area, our approach achieves par performance with other SOTA algorithms for detecting challenging pedestrians. In the nighttime scenario, as shown at the bottom of Figure 13, our proposed method successfully detects very distant dynamic objects in the highlighted dark-red-square area, demonstrating its superior long-range dynamic object perception capability.

### v2

@@ IV-E Performance Analysis on nuScenes-C Dataset
Method Clean IoU / mIoU Cam Crash IoU / mIoU Frame Lost IoU / mIoU Color Quant IoU / mIoU Motion Blur IoU / mIoU Bright IoU / mIoU Low Light IoU / mIoU Fog IoU / mIoU Snow IoU / mIoU OccCylindrical 44.94 / 28.67 41.76 / 23.38 40.84 / 21.87 43.64 / 25.38 41.45 / 22.57 44.78 / 27.88 43.48 / 22.17 44.59 / 27.38 38.92 / 18.29 DAOcc 45.00 / 30.50 42.75 / 22.35 42.16 / 20.78 43.88 / 27.46 42.75 / 26.25 43.72 / 26.78 41.25 / 21.15 43.92 / 26.87 40.72 / 19.42 TFusionOcc (T-P-25600) 45.49 / 30.04 42.59 / 22.34 41.97 / 20.78 43.14 / 24.42 40.45 / 20.04 44.81 / 29.03 41.24 / 18.62 44.14 / 27.43 40.58 / 15.62 TFusionOcc (T-SQ-25600) 46.14 / 30.88 42.59 / 23.50 41.93 / 22.04 43.03 / 25.65 38.36 / 20.25 45.41 / 29.92 41.99 / 20.98 44.90 / 28.85 40.09 / 16.33 TFusionOcc (T-SQ-IW-25600) 47.01 / 31.47 42.61 / 24.39 41.77 / 23.03 43.97 / 26.14 38.58 / 21.06 46.31 / 30.58 42.19 / 21.67 45.95 / 29.69 40.37 / 17.45
TABLE V: 3D semantic occupancy prediction results on nuScenes-C Camera Corruption benchmark. The best and second-best performance is indicated in bold red and bold blue, respectively.
In the camera corruption scenario, our model shows strong robustness against camera crashes, frame lost, brightness, and fog, as measured by the IoU and mIoU metric. Nevertheless, the motion blur and snow scenarios cause significant performance degradation in our model. With around 5.04% to 8.43% and 4.91% to 6.64% degradation under the IoU metric, 10.00% to 10.63% and 14.02% to 14.55% degradation under the mIoU metric for the motion blur and snow corruption case, respectively. This phenomenon could stem from the multi-stage feature fusion design strategy. Since frame lost and cam crash, both corruption results in the complete loss of information from the camera side. Our feature fusion module at each stage can easily filter out unnecessary visual features and rely solely on LiDAR features for occupancy prediction. As for brightness, fog, and low-light corruption, these conditions primarily degrade camera-side information quality and can be compensated for on the LiDAR side through our multi-stage feature fusion design. However, motion blur corruption causes structured information loss via convolution, bringing in inaccurate visual features into the feature fusion stage, resulting in a low-quality fused feature. Similarly, snow corruption introduces additional salt-and-pepper noise into the multi-stage feature fusion pipeline, resulting in a low-quality fused feature. Nevertheless, we consistently observe that more expressive primitives achieve stronger robustness under most corruption settings, following the trend T-SQ-IW >> T-SQ >> T-P, which further highlights the importance of geometric shape flexibility.
IV-F Performance Analysis on Varying Distance
@@
[…]
IV-G Qualitative Study
To demonstrate the effectiveness and robustness of our method, we compare TFusionOcc in three different T-Primitive settings, such as T-P-25600, T-SQ-25600 and T-SQ-IW-25600, against other SOTA algorithms in challenging daytime, rainy and nighttime scenarios, as shown in Fig. 10, with the primary difference area highlighted in each scenario.